OpenStudy (anonymous):

Solve by the substitution method: x=y-4 2x+2y=12

OpenStudy (anonymous):

2(y-4)+2y=12 2y-8+2y=12 4y-8=12 4y=20 y=5 to find x just plug y in to one of the other equations x=5-4 x=1 2(1)+2(5)=12? 2+10=12? 12=12!